# Machine learning surrogate models of many-body dispersion interactions in polymer melts

### v1

@@ 1 Introduction
While classical FF models are computationally efficient, they suffer from limited accuracy due to their inadequate description of polarization and many-body interactions. To overcome these drawbacks, machine learning force fields (MLFFs) [19, 20, 21, 22, 23] have emerged as promising solutions, offering quantum-level accuracy with competitive computational efficiency. MLFFs, which include a wide range of models from kernel-based methods to artificial neural networks (e.g., sGDML [24], SchNet [25, 26], PhysNet [27], GEMS [28], among others [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]), aim to learn the statistical relationship between chemical structures and the corresponding potential energy as well as atomic forces. Despite their advantages, many MLFFs are primarily trained for short-distance interactions and continue to rely on classical PW models for vdW interactions [42]. Some approaches, such as SpookyNet [43] and SO3LR [44], incorporate MBD-corrected reference data into their training while explicitly embedding parametric PW vdW functions within their architectures. These approaches, however, remain essentially parametric fittings to MBD and are limited in capturing many-body vdW effects beyond intramolecular interactions. To the best of our knowledge, there is currently no MLFF model explicitly designed as an MBD surrogate capable of capturing many-body vdW effects in large-scale systems.
@@
In this paper, we introduce a trimmed SchNet architecture that is specifically tailored for MBD force predictions in polymer melts. It is based on the SchNet architecture [25, 26], which we found an optimal choice for our problem due to its simplicity and superior adaptability (see also a more detailed discussion in Section 2.2). SchNet has demonstrated strong performance in molecular energy and force predictions, which is attributed to its radial basis function (rbf) encoding of structural information and continuous-filter convolutions. It remains scalable with respect to system size and ensures rotational invariance for energy prediction and rotational equivariance for force prediction. In our adaptation, we reduce the dimension and depth of the model while trimming the connection graph to retain only the most relevant interactions, which are then encoded using a reduced number of trainable rbfs. In addition, inspired by the repetitive structure of polymer chains, we introduce a unit-specific batching strategy that enhances training convergence. To validate the predictive capacity of the model, we apply it to three types of polymer melts: polyethylene (PE), polypropylene (PP), and polyvinyl chloride (PVC) and further assess its generalization capacity and robustness on mixed datasets and advanced MBD variants. A preliminary study is also conducted to evaluate temperature effects on the model’s performance. Furthermore, we demonstrate the physical interpretability of the model through analysis of the Hessian and discuss practical implementation of the model in molecular dynamics (MD) simulations, showcasing its applicability to large-scale systems. To increase the impact of this work, we open-source our codes and dataset, which are available in our repository [59].
@@ 2.1 Many-body dispersion method
$\boldsymbol{F}_{\boldsymbol{r}_{i}}^{\rm disp,MBD}=-\frac{1}{4}\mbox{Tr}\left[\boldsymbol{\Lambda}^{-1/2}\,\boldsymbol{S}^{\text{T}}\left(\nabla_{\boldsymbol{r}_{i}}\boldsymbol{C}^{\text{MBD}}\right)\boldsymbol{S}\right],\$ (4)
[…]
@@ 2.2 Surrogate modeling requirements for MBD in polymer melts
Even with conventional PW formulas, such as Eq. (1), computational feasibility is achieved by leveraging the decay properties of vdW dispersion to impose a cutoff range of interactions for each atom. This ensures that the overall computational complexity remains $O(N\times N_{\text{cut}})$, where $N_{\text{cut}}$ is the number of atoms within the cutoff neighborhood. To maintain efficiency, the MBD model must adopt a similar strategy. As illustrated in (a)-(c) of Fig. 1, MBD calculations are performed on the nearly spherical atomic clusters extracted from the full folded system, a strategy directly employed in the proposed surrogate modeling approach. Specifically, the surrogate model processes an atomic cluster as input and outputs the MBD force acting on its center atom, which is repeated for all atoms in the system. Moreover, accurately predicting MBD forces requires a significantly larger cutoff range compared to PW models, due to the inherently long-range nature of many-body correlations, as discussed in [11, 60]. This presents an additional challenge, as $N_{\text{cut}}$ scales cubically with the cutoff distance. On the positive side, the structural characteristics of polymer melts introduce notable regularity in the cutoff region. For a fixed $N_{\text{cut}}$, the extracted atomic clusters closely resemble spheres with similar radii. This structural uniformity benefits statistical ML surrogate modeling, as the limited number of atom types enables the model to focus primarily on encoding local geometric details rather than learning higher-level chemical features.
@@
To meet the above requirements, several well-established MLFF models, such as SchNet, PhysNet, and SpookyNet, are potential candidates due to their versatility and strong predictive performance. However, the size of these models is primarily constrained by their deep and complex architectures, which are designed for more generalized chemical systems, often incorporating enriched charge-dependent inputs or advanced attention-based modules [61]. Since computational efficiency is the most critical requirement for practical applications, we adopt a SchNet-based architecture, as it is the lightest among these models while offering robustness and greater adaptability. In the following section, we introduce our proposed model, the trimmed SchNet, and highlight key modifications tailored for the MBD surrogate problem.
@@ 2.3.2 Interactions by trimmed connections
In the application to MBD force prediction for polymers, we only need to calculate for the center atom in a given cutoff range. To maintain a fixed input dimension for the neural network, the atomic cluster is cut off by a fixed number of atoms $N=N_{\text{cut}}$, while the position tuple $\boldsymbol{R}$ is ordered by the atomic distance to the center atom. In the original version, a full SchNet leverages all $\sum^{N_{\text{cut}}}_{n=2}(n-1)$ connections between atoms, which restricts the cutoff range in practice. However, our specific goal, focusing only on the center atom, implies that a reduction of required connections is potentially feasible. In this paper, we implement a trimmed SchNet that remains limited connections, as shown in connection graph in Fig. 2, which updates the atom feature as follows:
@@
    Keep the full convolution to the atom $1$, i.e. the center atom:
@@
    Add $N_{\text{extra}}$ connections for the $p$ nearest neighbors of atom $1$:
@@
    The rest of the atoms only take reverse connections from the previous two updates:
@@
Trimming the peripheral connections effectively reduces the computational cost of the model, while the remaining extra connections to the nearest neighbors enhance the capability of the trimmed SchNet. The results presented in Section 3.2.1 highlight the importance of these additional connections for achieving highly accurate fitting. Moreover, in practice, a SchNet model can incorporate multiple interaction blocks in sequence, each with separate rbf encodings, allowing it to capture more complex and hierarchical features as the model depth increases. In this work, we opt for a single interaction block to balance accuracy with computational efficiency.
@@ 2.3.4 Force-based loss function
where $E$ here is only an energy-like representation after the projection since there will be no direct conservation for it. Nevertheless, the rbf encoding introduced above ensures rotationally invariance of $E$, so by construction, the force defined by Eq. (2) is rotationally equivariant.
With the single-force output, our trimmed SchNet is trained with a reduced loss function compared to the one for the original SchNet. If we denote this neural network as $h:(\boldsymbol{R},\boldsymbol{Z})\rightarrow\boldsymbol{F}_{\boldsymbol{r}_{1}}$ and its trainable parameters as $\boldsymbol{\theta}$, the mean squared error loss function used in this paper can be formulated as:
@@
In principle, a SchNet model that solely outputs energy is simpler and requires only half the computational cost [25]. However, our goal is to develop a surrogate model for MBD forces that can be directly integrated into MD simulations. If the model requires an additional step to compute forces, whether through a finite difference algorithm or through automatic differentiation (AD), it would inevitably introduce higher computational costs and pose additional challenges for deployment in external MD packages.
@@ 2.4 Three polymer melt systems: description and data generation
For training the neural network, we collect 60k data points for PE, 72k for PP, and 60k for PVC, with $10\%$ of each reserved for validation. These data points are extracted equally from different configurations for each type of polymer. Each data point is a $\left(\left(\boldsymbol{R},\boldsymbol{Z}\right),\hat{\boldsymbol{F}}_{\boldsymbol{r}_{1}}\right)$ tuple, representing an atomic cluster with $N_{\text{cut}}=1000$ atoms and the MBD force on its center atom (scaled by $10^{3}$ for numerical stability) generated by our open-source toolkit [45]. This number of $N_{\text{cut}}$ corresponds to a cutoff distance of approximately $14\,\text{\AA}$, significantly larger than the typical range of $6\text{-}9\,\text{\AA}$ used in classical PW methods due to the long-range nature of MBD. While $N_{\text{cut}}$ could be adjusted depending on the specific MBD variant and the density of the specific polymer melt, or reduced to accommodate stricter computational resource constraints, we select $N_{\text{cut}}=1000$ as a general-purpose value for demonstration, ensuring applicability across all cases in this paper. The selected clusters always include complete data points from the smallest repeating structural units in the polymers: methylene groups ($\text{C}\text{H}_{2}$) for PE, propylene monomers ($\text{C}_{3}\text{H}_{6}$) for PP, and vinyl chloride monomers ($\text{C}_{2}\text{H}_{3}\text{Cl}$) for PVC. This enables a special batching strategy to improve convergence of training as discussed in Section 3.2.3. Regarding the edge atoms of polymer chains, although they are not directly involved in the training due to their minor representation in the structure, the model demonstrates reasonable predictive capability for them after being trained on our sufficiently extensive dataset.
@@ 3.1 Model performance across polymers
Table 1: Performance ($\mathcal{E}_{\text{MARE}}$ (%)) of the model trained on single- and mixed-polymer datasets.
Train Test PE PP PVC
@@ C H all C H all C H Cl all
For single-polymer datasets, the model achieves low errors on the corresponding test datasets, with higher relative errors observed for carbon atoms. This is primarily due to their smaller force magnitudes and lower representation in the polymer structure compared to other atoms. Among the three polymers, the PVC case exhibits lower errors on carbon atoms, attributed to its distinct atomic composition and slightly higher force magnitudes of carbon atoms. However, these errors remain negligible for molecular simulations, given the larger mass of carbon atoms. Notably, compared to a PW model, the MBD model not only alters the magnitude of vdW dispersion forces but also significantly twists their directions due to many-body correlations. This directional alteration is crucial for accurately capturing many-body effects in the dynamical behavior of polymer melts, emphasizing the importance of force direction predictions alongside magnitudes. To assess this, we further examine the atomic force direction (angular) error $\mathcal{E}_{\text{angle}}\,(^{\circ})$, as defined in Eq. (16), for models trained on single-polymer datasets, corresponding to the bold results in Tab. 1. As shown in Fig. 5, force predictions for hydrogen atoms exhibit the highest angular accuracy across all three polymers, while chlorine atoms in PVC also maintain an angular error below $1^{\circ}$, in agreement with the observed $\mathcal{E}_{\text{MARE}}$. For carbon atoms, due to the aforementioned limitations, the angular error is generally higher, predominantly ranging between $1^{\circ}$ and $5^{\circ}$. Moreover, the model trained on PVC performs best on carbon atoms, aligning with the trend observed in the magnitude-based evaluation. Given the strong correlation and consistency between $\mathcal{E}_{\text{MARE}}$ and $\mathcal{E}_{\text{angle}}$, we do not present angular error analysis for the remaining cases.
Figure 6: Training convergence of the transfer learning experiment. Solid lines represent transfer learning (learning rate $10^{-4}$) between PE and PP: the PP-trained model fine-tuned on PE (blue) and the PE-trained model fine-tuned on PP (red). Dashed lines in corresponding colors indicate standalone training, which begins with a learning rate of $10^{-3}$ for the first 50 epochs and changes to $10^{-4}$ for the last 50 epochs.

### v2

@@ 1 Introduction
While classical FF models are computationally efficient, they suffer from limited accuracy due to their inadequate description of polarization and many-body interactions. To overcome these drawbacks, machine learning force fields (MLFFs) [55, 81, 59, 46, 67, 57] have emerged as promising solutions, offering quantum-level accuracy with competitive computational efficiency. MLFFs, which include a wide range of models from kernel-based methods to artificial neural networks (e.g., sGDML [16], SchNet [68, 69], PhysNet [82], GEMS [83], among others [7, 21, 89, 34, 42, 22, 19, 10, 11, 31, 30, 9, 8]), aim to learn the statistical relationship between chemical structures and the corresponding potential energy as well as atomic forces. Despite their advantages, many MLFFs are primarily trained for short-distance interactions and continue to rely on classical PW models for vdW interactions [4]. Some approaches, such as SpookyNet [80] and SO3LR [45], incorporate MBD-corrected reference data into their training while explicitly embedding parametric PW vdW functions within their architectures. These approaches, however, remain essentially parametric fittings to MBD and are limited in capturing many-body vdW effects beyond intramolecular interactions. To the best of our knowledge, there is currently no MLFF model explicitly designed as an MBD surrogate capable of capturing many-body vdW effects in large-scale systems.
@@
In this paper, we introduce a trimmed SchNet architecture that is specifically tailored for MBD force predictions in polymer melts. It is based on the SchNet architecture [68, 69], which we found an optimal choice for our problem due to its simplicity and superior adaptability (see also a more detailed discussion in Section 2.2). SchNet has demonstrated strong performance in molecular energy and force predictions, which is attributed to its radial basis function (rbf) encoding of structural information and continuous-filter convolutions. It remains scalable with respect to system size and ensures rotational invariance for energy prediction and rotational equivariance for force prediction. In our adaptation, we reduce the dimension and depth of the model while trimming the connection graph to retain only the most relevant interactions, which are then encoded using a reduced number of trainable rbfs. In addition, inspired by the repetitive structure of polymer chains, we introduce a unit-specific batching strategy that enhances training convergence. To validate the predictive capacity of the model, we apply it to three types of polymer melts: polyethylene (PE), polypropylene (PP), and polyvinyl chloride (PVC) and further assess its generalization capacity and robustness on mixed datasets and advanced MBD variants. A preliminary study is also conducted to evaluate temperature effects on the model’s performance. Furthermore, we demonstrate the physical interpretability of the model through analysis of the Hessian and discuss practical implementation of the model in molecular dynamics (MD) simulations together with a preliminary stability test, showcasing its applicability to large-scale systems. To increase the impact of this work, we open-source our codes and dataset, which are available in our repository [71].
@@ 2.1 Many-body dispersion method
$\boldsymbol{F}_{\boldsymbol{r}_{i}}^{\rm disp,MBD}=-\frac{1}{4}\mbox{Tr}\left[\boldsymbol{\Lambda}^{-1/2}\,\boldsymbol{S}^{\text{T}}\left(\nabla_{\boldsymbol{r}_{i}}\boldsymbol{C}^{\text{MBD}}\right)\boldsymbol{S}\right],\\$ (4)
where $\boldsymbol{\Lambda}_{lq}=\lambda_{l}\delta_{lq}$ is obtained from the diagonalization of $\boldsymbol{C}^{\text{MBD}}$ via a suitable SO(3NN) rotation matrix $\boldsymbol{S}$ that satisfies $\boldsymbol{\Lambda}=\boldsymbol{S}^{\text{T}}\boldsymbol{C}^{\text{MBD}}\boldsymbol{S}$. Summing over the eigenmodes reflects the collective contributions from the whole system, qualitatively differing from the pairwise force splitting in simpler PW models.
@@ 2.2 Surrogate modeling requirements for MBD in polymer melts
Even with conventional PW formulas, such as Eq. (1), computational feasibility is achieved by leveraging the decay properties of vdW dispersion to impose a cutoff range of interactions for each atom. This ensures that the overall computational complexity remains $O(N\times N_{\text{cut}})$, where $N_{\text{cut}}$ is the number of atoms within the cutoff neighborhood. To maintain efficiency, the MBD model must adopt a similar strategy. As illustrated in (a)-(c) of Fig. 1, MBD calculations are performed on the nearly spherical atomic clusters extracted from the full folded system, a strategy directly employed in the proposed surrogate modeling approach. Specifically, the surrogate model processes an atomic cluster as input and outputs the MBD force acting on its center atom, which is repeated for all atoms in the system. Moreover, accurately predicting MBD forces requires a significantly larger cutoff range compared to PW models, due to the inherently long-range nature of many-body correlations, as discussed in [74, 36, 2]. This presents an additional challenge, as $N_{\text{cut}}$ scales cubically with the cutoff distance. On the positive side, the structural characteristics of polymer melts introduce notable regularity in the cutoff region. For a fixed $N_{\text{cut}}$, the extracted atomic clusters closely resemble spheres with similar radii. This structural uniformity benefits statistical ML surrogate modeling, as the limited number of atom types enables the model to focus primarily on encoding local geometric details rather than learning higher-level chemical features.
@@
Several well-established MLFF models, such as SchNet, NequIP [9], MACE [8], SpookyNet [80], and SO3LR [45], satisfy these requirements and demonstrate strong general-purpose performance. However, many of these architectures are advanced derivatives of one another, designed for broad chemical spaces and enriched prediction targets, often incorporating higher-order features together with multiple layers of physically motivated interaction modules (e.g., attention-based modules [87]). While powerful, these additions introduce substantial computational overhead. In our case, the task is restricted to predicting MBD forces, a non-bonded interaction with a smoother landscape than the full interatomic physics, and is formulated for a single target atom per cutoff cluster. This narrower scope does not benefit from the additional complexity of such general-purpose models. Adapting them would likely involve removing most of their specialized components, effectively converging towards a model similar to SchNet. For this reason, we adopt a SchNet-based design, starting from a minimal yet robust baseline that avoids the overhead of unused functionality while retaining the required accuracy. This deliberate trade-off of generality for scalability makes the model well matched to our task. In the following section, we present our proposed trimmed SchNet and the dedicated modifications tailored to the MBD surrogate problem.
@@ 2.3.2 Interactions by trimmed connections
In the application to MBD force prediction for polymers, we only need to calculate for the center atom in a given cutoff range. To maintain a fixed input dimension for the neural network, the atomic cluster is cut off by a fixed number of atoms $N=N_{\text{cut}}$, while the position tuple $\boldsymbol{R}$ is ordered by the atomic distance to the center atom. In the original version, a full SchNet leverages all $\sum^{N_{\text{cut}}}_{n=2}(n-1)$ connections between atoms, which restricts the cutoff range in practice. However, our specific goal, focusing only on the center atom, implies that a reduction of required connections is potentially feasible. In this paper, we implement a trimmed SchNet that remains limited connections, as shown in connection graph in Fig. 2 (top left), which updates the atom feature as follows:
@@
    Keep the full convolution (orange lines) to the atom $1$, i.e. the center atom (red circle):
@@
    Add $N_{\text{extra}}$ connections (yellow lines) for the $p$ nearest neighbors (gray circles) of atom $1$:
@@
    The rest of the atoms (blue circles) only take reverse connections from the previous two updates:
@@
The trimmed connections, i.e., those disregarded from the original SchNet, are depicted as dashed gray lines. Trimming these peripheral connections effectively reduces the computational cost of the model, while the remaining extra connections to the nearest neighbors enhance the capability of the trimmed SchNet. The results presented in Section 3.2.1 highlight the importance of these additional connections for achieving highly accurate fitting. Moreover, in practice, a SchNet model can incorporate multiple interaction blocks in sequence, each with separate rbf encodings, allowing it to capture more complex and hierarchical features as the model depth increases. In this work, we opt for a single interaction block to balance accuracy with computational efficiency.
@@ 2.3.4 Force-based loss function
where $E$ here is only an energy-like latent quantity obtained after the projection layer of the network. Importantly, this quantity is not intended to represent a physically meaningful or conserved potential energy, as the model is trained on truncated atomic clusters extracted from a larger polymer melt system, for which a well-defined global energy does not exist. Nevertheless, the rbf encoding introduced above ensures rotationally invariance of $E$, so by construction, the force defined by Eq. (2) is rotationally equivariant.
With this force-only output, our trimmed SchNet is trained with a reduced loss function compared to the one for the original SchNet. If we denote this neural network as $h:(\boldsymbol{R},\boldsymbol{Z})\rightarrow\boldsymbol{F}_{\boldsymbol{r}_{1}}$ and its trainable parameters as $\boldsymbol{\theta}$, the mean squared error loss function used in this paper can be formulated as:
@@
In principle, a SchNet model that solely outputs energy is simpler and requires only half the computational cost [68]. However, our goal is to develop a surrogate model for MBD forces that can be directly integrated into MD simulations. If the model predicts energies only, forces must be obtained at every MD step via an additional differentiation operation, either through a finite difference algorithm or through automatic differentiation (AD). This extra step would inevitably introduce higher computational costs and pose additional challenges for deployment in external MD packages.
@@ 2.4 Three polymer melt systems: description and data generation
For training the neural network, we collect 60k data points for PE, 72k for PP, and 60k for PVC, with $10\%$ of each reserved for validation. These data points are extracted equally from different configurations for each type of polymer. Each data point is a $\left(\left(\boldsymbol{R},\boldsymbol{Z}\right),\hat{\boldsymbol{F}}_{\boldsymbol{r}_{1}}\right)$ tuple, representing an atomic cluster with $N_{\text{cut}}=1000$ atoms and the MBD force on its center atom (scaled by $10^{3}$ for numerical stability) generated by our open-source toolkit [70]. This number of $N_{\text{cut}}$ corresponds to a cutoff distance of approximately $14\,\text{\AA }$, significantly larger than the typical range of $6\text{-}9\,\text{\AA }$ used in classical PW methods due to the long-range nature of MBD. While $N_{\text{cut}}$ could be adjusted depending on the specific MBD variant and the density of the specific polymer melt, or reduced to accommodate stricter computational resource constraints, we select $N_{\text{cut}}=1000$ as a general-purpose value for demonstration, ensuring applicability across all cases in this paper. Our cutoff strategy is consistent with a recently reported statistical analysis of MBD forces in similar PE melt systems [74]. At this cutoff, the contributions from peripheral atoms decay smoothly towards negligible magnitudes, such that the outer region effectively serves as a buffer and changes in the cutoff neighborhood during MD lead to smooth force variations rather than discontinuous jumps, analogous to standard cutoff treatments in classical PW vdW models. The selected clusters always include complete data points from the smallest repeating structural units in the polymers: methylene groups ($\text{C}\text{H}_{2}$) for PE, propylene monomers ($\text{C}_{3}\text{H}_{6}$) for PP, and vinyl chloride monomers ($\text{C}_{2}\text{H}_{3}\text{Cl}$) for PVC. This enables a special batching strategy to improve convergence of training as discussed in Section 3.2.3. Regarding the edge atoms of polymer chains, although they are not directly involved in the training due to their minor representation in the structure, the model demonstrates reasonable predictive capability for them after being trained on our sufficiently extensive dataset.
@@ 3.1 Model performance across polymers
Table 1: Performance ($\mathcal{E}_{\text{MARE}}$ (%)) of the model trained on single- and mixed-polymer datasets. Values in parentheses indicate standard deviations of $\mathcal{E}_{\text{MARE}}$ (%).
Train Test PE $P$ PVC
@@ C H all C H all C H Cl all
For single-polymer datasets, the model achieves low errors on the corresponding test datasets, with higher relative errors observed for carbon atoms. This is primarily due to their smaller force magnitudes and lower representation in the polymer structure compared to other atoms. Among the three polymers, the PVC case exhibits lower errors on carbon atoms, attributed to its distinct atomic composition and slightly higher force magnitudes of carbon atoms. However, these errors remain negligible for molecular simulations, given the larger mass of carbon atoms. Notably, compared to a PW model, the MBD model not only alters the magnitude of vdW dispersion forces but also significantly twists their directions due to many-body correlations [74]. This directional alteration is crucial for accurately capturing many-body effects in the dynamical behavior of polymer melts, emphasizing the importance of force direction predictions alongside magnitudes. To assess this, we further examine the atomic force direction (angular) error $\mathcal{E}_{\text{angle}}\,(^{\circ})$, as defined in Eq. (16), for models trained on single-polymer datasets, corresponding to the bold results in Tab. 1. As shown in Fig. 5, force predictions for hydrogen atoms exhibit the highest angular accuracy across all three polymers, while chlorine atoms in PVC also maintain an angular error below $1^{\circ}$, in agreement with the observed $\mathcal{E}_{\text{MARE}}$. For carbon atoms, due to the aforementioned limitations, the angular error is generally higher, predominantly ranging between $1^{\circ}$ and $5^{\circ}$. Moreover, the model trained on PVC performs best on carbon atoms, aligning with the trend observed in the magnitude-based evaluation. Given the strong correlation and consistency between $\mathcal{E}_{\text{MARE}}$ and $\mathcal{E}_{\text{angle}}$, we do not present angular error analysis for the remaining cases.
Figure 6: Training convergence of the transfer learning experiment. Solid lines represent transfer learning (learning rate $10^{-4}$) between PE and $P$: the PP-trained model fine-tuned on PE (blue) and the PE-trained model fine-tuned on PP (red). Dashed lines in corresponding colors indicate standalone training, which begins with a learning rate of $10^{-3}$ for the first 50 epochs and changes to $10^{-4}$ for the last 50 epochs.
@@
In addition to the mean error values, Tab. 1 includes their standard deviations in parentheses. In most cases, the standard deviations are comparable to or larger than the mean values due to the skewed and heavy-tailed nature of the error distributions, as confirmed by the histograms presented in Fig. 5 in this section and Fig. 9 in Section 3.3. Such distributions reflect the intrinsic complexity of the MBD force landscape and configurational diversity in polymer melt systems. For clarity, standard deviations are omitted in subsequent tables, as they focus on model comparisons and the overall statistical trends remain consistent.